# MMUTF: Multimodal Multimedia Event Argument Extraction with Unified Template Filling

### v1

@@ A.3 Threshold Tuning
Transaction:Transfer-Money [Giver] paid Recipient] with [Money] for the benefit of [Beneficiary] at [Place].
@@
A.4 Implementation Details
In our experiments, we use the implementation of the Transformers (Wolf et al., 2020) (v4.41.0) library in conjunction with PyTorch (v2.3.0). For all runs, we evaluated the performance at each epoch on the ACE2005 test set and selected the best-performing checkpoint. Unless otherwise mentioned, we use T5-base [^3] https://huggingface.co/google-t5/t5-base (Raffel et al., 2020) with  [^2] M parameters as text encoder-decoder and CLIP-base [^4] https://huggingface.co/openai/clip-vit-base-patch16 (Radford et al., 2021) with 85M parameters as vision encoder model. All models are loaded in FP16 and executed on A100 GPUs with 40GB HBM using the same compute node running CUDA 12.3 and NVIDIA device driver version 545. The average training time for each run did not exceed 3h.
A.5 Event Templates

### v2

@@ A.2 Threshold Tuning
Transaction:Transfer-Money [Giver] paid [Recipient] with [Money] for the benefit of [Beneficiary] at [Place].
@@
Appointing [Selector] appoint to [Body] [Official] as [Role] [Function] in [Place].
Arranging [Agent] arrange [Theme] in [Configuration] with [Instrument].
Fining [Speaker] fine [Payer] [Fine] [Reason] with [Instrument] in [Place].
@@
A.3 Implementation Details
In our experiments, we use the implementation of the Transformers (Wolf et al., 2020) (v4.41.0) library in conjunction with PyTorch (v2.3.0). For all runs, we evaluated the performance at each epoch on the ACE2005 test set and selected the best-performing checkpoint. Unless otherwise mentioned, we use T5-base (Raffel et al., 2020) with  [^2] M parameters as text encoder-decoder and CLIP-base (Radford et al., 2021) with 85M parameters as vision encoder model. All models are trained with mixed precision and executed on A100 GPUs with 40GB HBM using the same compute node running CUDA 12.3 and NVIDIA device driver version 545. The average training time for each run did not exceed 3h.
Model Textual F1 Visual F1 MM F1
BERT ViT 37.9 19.2 26.6
CLIP 37.6 19.7 26.6
Data2Vec 38.4 18.3 26.1
BART ViT 38.1 19.2 24.9
CLIP 38.5 20.4 27.9
Data2Vec 39.0 18.4 27.0
T5 ViT 38.3 19.5 27.6
CLIP 38.2 20.9 27.4
Data2Vec 38.5 18.5 27.0
Table 6: Results for text and vision models.
A.4 Event Templates
@@
A.5 Backbone Models
We compare different text and vision model configurations in Table 6. For the pre-trained text models, we compare the following encoder and encoder-decoder variants: BERT [^3] https://huggingface.co/google-bert/bert-base-uncased (Devlin et al., 2019), BART [^4] https://huggingface.co/facebook/bart-base (Lewis et al., 2020), and T5555https://huggingface.co/google-t5/t5-base (Raffel et al., 2020). Regarding BERT, we use two distinct models for candidate and query representations and add cross-attention to the query encoder. For the pre-trained vision models, we rely on ViT [^6] https://huggingface.co/google/vit-base-patch16-224 (Dosovitskiy et al., 2021), CLIP [^7] https://huggingface.co/openai/clip-vit-base-patch16 (Radford et al., 2021), and Data2Vec [^8] https://huggingface.co/facebook/data2vec-vision-base (Baevski et al., 2022). All models represent their corresponding base variants, as we did not observe substantial improvements with larger model sizes.
Overall, the results demonstrate robust performance across different model configurations. However, we observe that CLIP generally outperforms the other vision models in terms of visual and multimedia F1 scores. We hypothesize that pre-training on large image-caption datasets is beneficial for the M2E2 task.
Template Textual F1 Visual F1 MM F1
w/ event type
Concatenation 38.0 18.9 27.9
Standard 38.3 17.9 25.3
Enriched 38.5 16.8 25.2
w/o event type
Concatenation 38.7 19.0 28.0
Standard 38.2 20.9 27.4
Enriched 38.0 19.7 28.1
Table 7: Results for different prompt templates.
A.6 Prompt Analysis
In Table 7, we investigate prompt template variations and conduct experiments with and without event type prefixes. For event type prefixes, we prepend the event type name to the event template (e.g., "Conflict Attack: <Event Template>"). We compare three different prompt variations: Concatenation simply concatenates the argument role names from the event templates. Standard represents the manually crafted event templates used in this work (cf. Table 5). Enriched adds argument role definitions after each argument role name.
The results show that event templates without event type prefixes generally outperform their counterparts across all variations. We conjecture this is because similar arguments appear in multiple events, and event type prefixes may impede the transfer of knowledge between roles. All variations perform similarly for textual EAE, while event type prefixes slightly impair model performance for visual and multimedia EAE.
A.7 Error Analysis
While our method achieves significant improvements across all SOTA models for the textual EAE task, we still recognize shortcomings for the visual modality. Therefore, we analyse the impact of visual event mentions and visual candidates predicted by YOLOv8 (Jocher et al., 2023). With gold triggers (i.e., annotated visual event mentions), we achieve an absolute improvement of 16% in visual F1 which highlights the importance of reliable ED. However, the inclusion of gold candidates (i.e., annotated bounding boxes) demonstrates an absolute improvement of 35.5% visual F1 which further enhances the gold triggers by 19.5%. These results suggests that the correct selection of candidates is crucial and the joint modeling of candidate generation and EAE poses a promising direction.